Be sure to answer all parts. What is the maximum mass (in grams) of each of the following soluble salts that can be added to 1.0 x 102 mL of 0.050 M Bac12 without causing a precipitate to form? (Assume that the addition of solid causes no change in volume.) (a) (NH402 SO4 X 10 g (Enter your answer in scientific notation.) (b) Pb(NO3) (c) NaF
Answer #1

The stoichiometry of the reactants to be 1.

First BaCl2,

Molecular Wt. of BaCl2 = 208.23 g/mol

Using the equatio of Molarity

Molarity = No of Moles / Volume in Litres = (Wt. / Molecular Wt.) / Volume in Litres

So, No. of moles of of BaCl2 = 0.050 / 10 = 0.005 moles (limiting reagent)

So all other reactants will be at max of 0.005 moles.

Now,

Wt. of (NH4)2SO4 = 0.005 * 132.14 = 0.661 gram = 66.1 X 10-2 g (Mol Wt of (NH4)2SO4 = 132.14 g/mol)

Wt. of Pb(NO3)2 = 0.005 * 331.2 = 1.66 gram   (Mol Wt of Pb(NO3)2 = Pb(NO3)2 g/mol )

Wt. of NaF = 0.005 * 41.988 = 0.21g (Mol Wt of NaF = 41.988 g/mol )

Fell free to comment.

Thank You